Road Warrior Corporation began operations early in the current year, building luxury motor homes. During the year, the company started and completed 50 motor homes at a cost of $60,000 per unit. Of these, 48 were sold for $100,000 each and two remain in finished goods inventory. In addition, the company had six partially completed units in its factory at year-end. Total costs for the year (summarized alphabetically) were as follows.

Direct materials used $ 765,000
Direct labor 821,000
Income tax expense 100,000
General and administrative expenses 500,000
Manufacturing overhead 1,690,000
Selling expenses 500,000

Required:

a. Compute the total manufacturing costs charged to work in process for the current year.

b. Compute the cost of finished goods manufactured for the current year.

c. Compute the cost of goods sold for the current year.

d. Compute the gross profit on sales for the current year.

e. Compute the ending inventories of (1) work in process and (2) finished goods for the current year.
Answer #1

For calculating the amounts, we will make use of a cost sheet, but we will be using it in the reverse to find the answers. First, let's look at the completed cost sheet.

No. of units Amount
Direct Materials        765,000
+Direct Labor        821,000
+Direct Expenses nil
Prime Cost 56     1,586,000
+ Manufacturing Overhead     1,690,000
cost upto point     3,276,000
Add: opening work in progress nil
Less: closing work in progress 6        776,000
Cost of Manufacturing 50     2,500,000
+ Administrative Expenses        500,000
Cost of Production 50     3,000,000
Add: opening finished good inventory nil
Less: closing finished goods inventory 2        120,000
Cost of Production of goods sold 48     2,880,000
+ Selling Expenses        500,000
Cost of Goods Sold 48     3,380,000
Gross Profit     1,420,000
Sales 48     4,800,000

Let us analyze the cost sheet to understand how we derived the numbers.

Note: certain formulas are followed in alphabets in brackets are answers for the particular question that the alphabet represents.

In such questions, make a cost sheet and fill in the numbers that are easily derived at. From there, it will be easier to calculate the other amounts asked for. In the cost sheet I have made above, I added a column for no on units for which the particular cost is incurred, so that it will be easier for you to understand the same.

50 units were started and completed in the year at $60,000 cost for each unit. This is the Cost of Production, The total of cost of production incurred is 50 x $60,000 = $ 3,000,000.

50 units were started and completed in the year, while 6 still remain partially completed, which means, they remain in work in progress. All costs up to manufacturing overhead have been incurred by all of these 56 units. We need to segregate the work in progress cost from the cost of completed units. We can calculate the cost of the completed units using the below formulas:

Cost of production = Cost of manufacturing + administrative expenses

Therefore, cost of manufacturing = cost of production - administrative expenses

= 3,000,000 - 500,000 = $ 2,500,000 (b.)

So, the 50 units incur $ 2,500,000 for manufacturing. Any costs incurred in the manufacturing process over this amount has been incurred by the goods which still remains incomplete. The total cost of manufacturing incurred in the company for the 56 units is calculated below:

Direct materials + Direct Labor + Manufacturing Overhead = 765,000 + 821,000 + 1,690,000

= $ 3,276,000

As we calculated, the cost of manufacturing 50 units is only $ 2,500,000. Thus the cost of manufacturing the work-in-progress is the difference between the total manufacturing cost of 56 units and manufacturing cost of 50 units.

= 3,276,000 - 2,500,000 = $ 776,000 (a.)

Out of the 50 units completed, only 48 units were sold that year, while 2 still remained in the factory. Hence, any selling cost incurred will only be for the 48 units sold, and not for all 50 units produced. This is why $60,000 is the cost of producing each unit and not the cost of goods sold.

Cost of good sold = cost of production of goods sold + selling expenses

48 units were sold that year. The cost of production for these 48 units are 48 x $ 60,000 = $ 2,880,000. The selling expenses incurred is $ 500,000.

Therefore, cost of goods sold = 2,880,000 + 500,000 = $ 3,380,000 (c.)

Sale price of each unit is $ 100,000. Therefore, total sales value is 48 x $ 100,000 = $ 4,800,000.

Gross profit = Sales price - Cost of goods sold

= 4,800,000 - 3,880,000 = $ 1,420,000 (d.)

Let us now answer the questions asked:

a. Manufacturing cost charged to work in process = $ 776,000

This is the cost incurred by the 6 units which are partially completed. It is the excess we calculated over the cost of manufacturing the 50 units.

b. Cost of finished goods manufactured for the year = $ 2,500,000

note: the question is about the cost of finished goods manufactured, which probably means the manufacturing cost which as per calculations is $ 2,500,000. In case the question was the cost of producing the finished good, it is $ 3,000,000. Please, confirm what the question specifically asks for before you submit your answer.

c. Cost of goods sold for the year = $ 3,380,000

This is the cost incurred for the 48 units that were sold in the year.

d. Gross profit = $ 1,420,000

The difference between the sales price for the 48 units and the total cost incurred for the same including selling costs.  

e. 1. Ending inventory of work in process = 6

this is the units that are still work in process or are only partially completed. As mentioned in the question, it is 6 units that are partially completed.

2. Ending inventory of finished goods = 2

this is the number of units out of finished goods that are not sold and still remain in inventory. This is given by 50 - 48 =2 units.